Seven apples and one banana cost 98 cents. Two apples and three bananas cost 66 cents. Findthe cost of one apple and the cost of one banana.
Answer by tinbar(133)

let a = apples and b = banana
then we have:
1) (7*a) + (1*b) = $0.98 and,
2) (2*a) + (3*b) = $0.66
now from statement 1) let's re arrange and isolate b, we get that b = $0.98-(7*a)
now since the apples and bananas are the same things in both 1) and 2) we can take our rearrangement of b in 1) and replace it in the b for 2)
so we'll get, in 2), that (2*a) + (3*(0.98-7a)) = 0.66
you should be able to do the rest on your own, from the rearrangement in the second statement you only have one variable, the apples, collect like terms and solve for a. once you do that, plug the value of a in either 1) or 2) and solve for b, you'll get the same answer regardless of choosing 1) or 2)